OpenStudy (anonymous):

What is the relationship between the lines determined by the following two equations? 6x – 2y = 16 y = 3x – 8 A. neither parallel nor perpendicular B. they are the same line C. parallel D. perpendicular I think its d

OpenStudy (anonymous):

@Auxuris

OpenStudy (anonymous):

-2y=16-6x -y=8-3x y=3x-8 answer is D as they have the same gradient and constant, they are the same line^^

OpenStudy (anonymous):

eeek sorry i meant B not D /shot same line, B

OpenStudy (anonymous):

so first they are C and now their B

OpenStudy (anonymous):

Uh no they're B. Always B.

OpenStudy (anonymous):

As i said i'm a careless-mistake person /hides